# Theorem-Carrying-Transaction: Runtime Certification to Ensure Safety for Smart Contract Transactions

### v1

@@ 3.2 Theorem to prove safety of a transaction
The notion of theorem is important in TCT. We explain it using the code paths in the two attacks. For Attack 1, the code path of the transaction is shown in 1, using the Boogie language [11]. A require check in Solidity corresponds to an assume in Boogie, because passing a require check means that the code path is subject to the constraint from the symbolic reasoning perspective. The operators + and - are modeled as functions evmadd and evmsub, correspond to opcodes ADD and SUB in the Ethereum yellow paper [17].
@@
Code 1: Code path of Attack 1 in the Boogie language.
2 lists some axioms in TCT’s basic library. They are not specific to the discussed case. In the axioms, m is a map from address to int. Notation m[a:=v] denotes a map that is almost identical to m, but its index a contains the value v.
function sum(m: [address] int) returns (int);
axiom forall m: [address] int, a:address ,v:int::
sum(m[a:=v]) == sum(m) - m[a] + v;
axiom forall m: [address] int ::
((forall a:address :: 0 <= m[a]) ==>
(forall a:address :: m[a] <= sum(m)));
function evmadd(a,b:int) returns (int);
axiom forall a,b:uint256 ::
a+b < 2^256 && a+b >= 0 ==> evmadd(a,b) == a+b;
axiom forall a,b: uint256 ::
a+b >= 2^256 ==> evmadd(a,b) == a+b-2^256;
function sub(a,b:int) returns (int);
axiom forall a,b: uint256 ::
a-b < 2^256 && a >= b ==> evmsub(a,b) == a-b;
axiom forall a,b: uint256 ::
a < b ==> evmsub(a,b) == a-b+2^256;
Code 2: Axioms of mathematical operations.
Because Prop. 1 is specified as invariants of ERC20, the derived MultiVulnToken contract should obey them. Invariants are conditions every execution should maintain, i.e., assuming the invariants are true in the beginning, they should be true in the end. 3 represents the proof obligation of the code path in 1. The first two inv lines assume the invariants, and the last two inv lines assert them. The hyp line (i.e., hypothesis) specifies a condition $\varphi(v,s)$, where $v$ represents parameters of the transaction and $s$ represents the blockchain storage states. The hypothesis is the condition under which the invariants will hold.
hyp: assume $\varphi$(v,s);
@@
Code 3: Proof obligation of the code path in 1.
Suppose $\varphi(v,s)$ is (0 $\leq$ totalSupply < $\texttt{2}^{\texttt{255}}$) $\land$ (0 $\leq$ _value < $\texttt{2}^{\texttt{255}}$) $\land$ (0 $\leq$ _fee < $\texttt{2}^{\texttt{255}}$), which consists of the contract variable totalSupply in $s$ and the parameters _value and _fee in $v$ of the entry function transferProxy. With the hypothesis, Boogie can successfully verify 3, meaning that the theorem $\tau$ in 4 is valid.
@@
$\varphi$(v,s) = (0 $\leq$ totalSupply < $\texttt{2}^{\texttt{255}}$) $\land$
(0 $\leq$ _value < $\texttt{2}^{\texttt{255}}$) $\land$ (0 $\leq$ _fee < $\texttt{2}^{\texttt{255}}$),
ph = 0xf222eb3c699342dcf8f4f64a86ba0862e
29fc8d5d28e21940ee0b32cf3eb06ca
@@
In the theorem $\tau$, the first part $f$ specifies the entry function of the transaction that the theorem pertains to. For readability of this paper, we use the name multiVulnToken::transferProxy, but in reality, the contract is identified by its address, and the function is identified by a 4-byte hexical value called “function selector” [18]. The second part is the hypothesis $\varphi(v,s)$. The third part is the path hash, which is a Keccak-256 hash identifying an EVM code trace. Section 5.4 will describe how it is computed in our current prototype. In summary, the general form of a theorem is $\tau:=(f,\varphi(v,s),ph)$, which means: for every transaction invoked by calling $f$ when $\varphi$ is satisfied, if it is completed (i.e., not reverted by EVM) and its path hash equals $ph$, then all the assertions along the code path are guaranteed to hold. Obviously, this theorem is not applicable to the transaction in Attack 1 because it violates (0 $\leq$ _value < $\texttt{2}^{\texttt{255}}$) $\land$ (0 $\leq$ _fee < $\texttt{2}^{\texttt{255}}$).
The generality of the theorem can be understood from three aspects. First, the hypothesis not only specifies the conditions for the three variables, but also implicitly states that nothing else matters, such as other parameters in the call or other storage variables. For example, addresses msg.sender, _from and _to can be arbitrarily aliased to each other, and the assertions will still hold. The theorem is thus a powerful statement. Second, the generality is far enough for any instance of the token contract in reality, as it is hard to imagine why a cryptocurrency needs a $2^{255}$ total supply (equivalent to $7*10^{66}$ per person in the world). Unlike static code verifiers, which hinge on the ability to expose unexpected boundary conditions, TCT enables programmers to focus on conditions only within the boundary of their intentions. Third, we will explain that TCT utilizes generality of the hypotheses to reuse theorems. It is crucial for TCT’s near-zero runtime overhead.
hyp: assume $\varphi$(v,s);
@@
Theorem about the reentrant code path. We now do a similar exercise about a transaction entering via clear(...). If it does not exploit the reentrancy pattern (i.e. if transition D in Figure 2 becomes a return), then a hypothesis like $\varphi(v,s)$ := (0  $\leq$  totalSupply < $\texttt{2}^{\texttt{255}}$) $\land$ (_to $\neq$ msg.sender) is sufficient to ensure the invariants. However, if the transaction exploits the reentrancy pattern as in Attack 2, the code path is different, as shown in 5. The two lines marked with “*” are repeated 9 more times. Hence the theorem for the non-reentrancy transaction is not appliable to the reentrancy one, because of the path hash mismatch. The programmer is unlikely to anticipate the reentrancy code path before the year 2016, but it is fine because security of TCT does not require the programmer to anticipate every non-intention. Since it is the attacker who discovers this unintended code path, it is the attacker’s obligation to prove the corresponding theorem. The aforementioned $\varphi(v,s)$ will fail the proof. The attacker can make the hypothesis $\varphi(v,s)$ := (balances[msg.sender] = 0) to prove the theorem, but it makes the attack harmless as the transaction is basically a no-op repeated 10 times.
4 The TCT Protocol
The TCT protocol requires only small extensions from the current Ethereum system. The protocol consists of three parties as shown in Figure 3 – the transaction issuer (e.g., MetaMask [19]), a Web-API service (e.g., infura.io) and the Ethereum network (i.e., the blockchain). In today’s Ethereum ecosystem, the Web-API service provides various functionalities, including gas-estimate for a transaction and failure warnings. It consists of a Web-API interface and a node connecting to the Ethereum network. In TCT, the Web-API service provides additional functionalities for storing and retrieving theorems. The most unique component for TCT is the theorem repository on each node, which contains the theorems that have been proven.
@@ 4 The TCT Protocol
Trusted computing base. TCT’s trusted computing base (TCB) is entirely at the infrastructure level. It does not include any contract-level assumption (e.g. about the reference-topology, contracts of interest, contracts that have been formally verified, etc), because TCT faithfully treats the entire GP as the target program. Our TCB includes Ethereum’s TCB and TCT-additional TCB. The former includes the Ethereum protocol, EVM, and the verified-compilation to attest a contract’s EVM code to its claimed source code [20]. The latter includes the TCT protocol (in this section), the Boogie verifier and a module that translates a trace from EVM to Boogie (in Section 5).
Protocol flow A. The first protocol flow is shown in Figure 3, which describes how the transaction issuer submits a transaction TX to the Ethereum network. First, TX is submitted to the Web-API service. In addition to the existing functionalities (e.g., gas-estimation, etc.), the service tries to find a theorem applicable to TX. Applicability means that the parameters ($v$) of TX and the storage states ($s$) satisfy the hypothesis $\varphi(v,s)$ of the theorem, and the path hash of executing TX equals the one in the theorem. Suppose such a theorem $\tau$ is found, the Web-API service sends $\tau$ to the transaction issuer, who then submits the tuple (TX, $\tau$) to the blockchain. Every blockchain node ensures the existence of $\tau$ in the repository, checks $\varphi(v,s)$, then executes TX. When the execution is completed, the actual path hash is compared against the one in theorem $\tau$. If they are not equal, the execution is reverted.
Like the current Ethereum protocol, the TCT protocol does not strictly require the interaction with the Web-API service before submitting a transaction. If the issuer already knows $\tau$, the protocol flow can start from “Send(TX,$\tau$)”.
Figure 3: Protocol flow A – when a transaction has an applicable theorem to prove its safety.
Protocol flow B. The flow in Figure 4 is used for any person, likely a programmer, to add a theorem about a known code trace into the theorem repository. CT represents the EVM bytecode sequence of the code trace. The issuer submits tuple (CT, $\tau$) to the blockchain. Each node computes the path hash of CT, and checks it against the one in the theorem. If they do not match, the theorem is rejected. Otherwise, the protocol proceeds. Section 5 will describe our technology that converts CT and $\tau$ into a Boogie code trace. The Boogie verifier is used to prove the code trace. If successful, the theorem is accepted into the repository. Note that the verification step burns a certain amount of gas to prevent denial-of-service attacks.
The goal of protocol flow B is for a programmer to pre-populate the repo with all anticipated theorems when the contract is deployed. The code traces can be recorded when the contract is tested on a local EVM or other test environments.
Figure 4: Protocol flow B – adding a theorem about a code trace into the repository.
Protocol flow C. In protocol flow A, an applicable theorem is found for TX. If no such theorem is found, the steps are shown as protocol flow C in Figure 5, which combines flow A and flow B. When the node receives $(TX,\tau)$, it test-runs TX to get the EVM code trace CT, then follows the same steps in flow B to prove and accept $\tau$ into the theorem repo. The remaining steps are the same as in flow A, which execute TX for real.
Figure 5: Protocol flow C – when a transaction has no applicable theorem to prove its safety.
Expectation about the three flows. We expect that flow A is the primary flow to process real users’ transactions, because most transactions are expected to be normal and have applicable theorems in the repository. Most theorems are accepted via flow B before the contracts process real user’s transactions, because the programmers are supposed to run test cases that cover all normal scenarios they can imagine. Flow C happens only when a real user issues a transaction that goes through a code path unseen in all test cases, which is likely an attack, e.g., the reentrancy attack in Section 3. However, if the transaction is legitimate, the user needs to take the burden to provide a new theorem, which should be very rare for a well-tested contract. Even in this case, the only real burden for the user is to provide the hypothesis $\varphi(v,s)$ for the theorem. Proving a straight-line code trace is easy. Our experience with Boogie is that all traces we have encountered are proven automatically.
@@
Optimizations for low storage and communication overheads. To store theorems efficiently in the repository, all theorems that share the same entry function and hypothesis but differ in their path hashes can be combined as a multi-path theorem, which states that, if the hypothesis is satisfied, then any code path in this multi-path set is safe. With this design, the number of theorems is the number of hypotheses, regardless of the number of possible code paths under each hypothesis. The amortized cost to store a new code path is only one Keccak-256 value, i.e., 32 bytes.
The instructions in a code path do not need to be stored, because they are only used during the theorem-proving time. Once it is done, the protocol no longer needs the exact EVM instructions, so the path hash suffices. Once a theorem is stored in the repository, a Keccak-256 hash can be computed for it. We call it theorem hash. In all network messages, $\tau$ can be represented by its theorem hash. Because most transactions go through flow A, we can conclude that the network overhead of TCT is 64 bytes per transaction, due to the messages “($\tau$)” and “Send(TX,$\tau$)”.
5 Prototype
In this section, we describe a prototype of TCT that has implemented all the essential mechanisms. The prototype is tested using real-world ERC20 contracts and a sophisticated application – Uniswap-V2, which we will describe in Section 6. These case-studies demonstrate that TCT is applicable to complex real-world multi-contract transactions, proving design-level properties that are difficult for traditional verification approaches.
5.1 Specifying properties in Solidity
The properties are specified using the NatSpec notation of Solidity [10]. A contract invariant is specified using the tag @custom:invariant above the contract’s definition, and a function postcondition is specified using the tag @custom:postcondition above the function’s definition. For example, 6 specifies the invariants for the ERC20 contract. The invariants could also be specified for the MultiVulnToken contract, but we specify them for ERC20 to show that properties of a base class are ensured by TCT for all derived classes, which demonstrates behavioral subtyping.
///@custom:invariant sum(this.balances) == this.totalSupply
contract ERC20 {
address public owner;
uint256 public totalSupply
mapping(address => uint256) balances;
Function declarations such as transfer(...), allowance(...), transferFrom(...), ...
Code 6: Specifying the invariants for ERC20.
5.2 Proving a theorem for a transaction
@@
Figure 6 shows how a theorem about a transaction is proven in the prototype system. We use Remix IDE [21] as the transaction issuer. The node consists of a Geth executable and other components. From Remix, we deploy a contract and issue a transaction. Geth returns a TX-hash that identifies the transaction. Note that TX-hash is an Ethereum concept, not to be confused with our TCT concept path hash. Based on the debug_traceTransaction feature of Geth, we build a trace fetcher, which uses TX-hash to get the EVM code trace.
Figure 6: Proving a theorem for a concrete code path.
@@
The main component we build is the concolic executor, which will be described next. The execution result of it is a straight-line Boogie code like the one in 3. The input includes the EVM code trace, the theorem, and Solidity compilation symbol information. The properties specified in the NatSpec notation are in the symbol information. When TCT is deployed, the symbol information should come from a verified-compilation service, such as the Etherscan service named Verify & Publish Contract Source Code [20]. The service attests that a contract’s EVM code is compiled from the claimed source code by an official Solidity compiler.
5.3 Concolic executor
The concolic executor is basically an EVM, which sequentially executes the instructions in a code trace. The main difference from a classical EVM is that the computation of each instruction is performed symbolically. The data structure that we define to hold a symbolic value is named symbolic value tree, or SVT.

### v2

@@ 3.2. Theorem to prove safety of a transaction
The notion of theorem is important in TCT. We explain it using the code paths in the two attacks. For Attack 1, the code path of the transaction is shown in LABEL:code:attack1 using the Boogie language (leino_boogie_2, 13). A require check in Solidity corresponds to an assume in Boogie. The operators + and - are modeled as functions evmadd and evmsub, correspond to opcodes ADD and SUB. TCT has a Boogie library to define axioms about these opcodes based on the Ethereum yellow paper (wood2014ethereum, 39). Axioms relevant to LABEL:code:attack1 are shown in Appendix A.1.
@@
Code 2: Code path of Attack 1 in the Boogie language.
Let’s assume Property 1 is specified for ERC20 as in LABEL:code:inv. TCT ensures behavioral subtyping (liskov1994behavioral, 15), so the derived MultiVulnToken contract must obey all properties of its base classes. LABEL:code:proof represents the proof obligation in this case. The first two inv lines assume the invariants, and the last two inv lines assert them. The grd line (i.e., “guard”) specifies a condition $\varphi(s,f.p)$, where $s$ represents the blockchain storage states and $f.p$ represents parameters of the entry function $f$. The guard is a condition under which the invariants will hold.
grd: assume $\varphi(s,f.p)$;
@@
Code 3: Proof obligation of the code path in LABEL:code:attack1.
Suppose $\varphi(s,f.p)$ is (0 $\leq$ totalSupply < $\texttt{2}^{\texttt{255}}$) $\land$ (0 $\leq$ _value < $\texttt{2}^{\texttt{255}}$) $\land$ (0 $\leq$ _fee < $\texttt{2}^{\texttt{255}}$), which consists of the contract variable totalSupply in $s$ and the parameters _value and _fee in $f.p$ of the entry function transferProxy. With the guard, Boogie successfully verifies LABEL:code:proof, i.e. theorem $\tau$ in LABEL:code:exmpl1 is valid.
@@
$\varphi(s,f.p)$ = (0 $\leq$ totalSupply < $\texttt{2}^{\texttt{255}}$) $\land$ (0 $\leq$ _value <
$\texttt{2}^{\texttt{255}}$) $\land$ (0 $\leq$ _fee < $\texttt{2}^{\texttt{255}}$),
ph = 0xf222eb3c699342dcf8f4f64a86ba0862e29fc8d5d2
8e21940ee0b32cf3eb06ca
@@
In theorem $\tau$, the first part $f$ specifies the entry function of the transaction that the theorem pertains to. For readability of this paper, we use the name multiVulnToken::transferProxy, but in reality, the contract is identified by its address, and the function is identified by a 4-byte hexical value called “function selector” (soliditydoc, 30). The second part is the guard $\varphi(s,f.p)$. The third part is the path hash, which is a Keccak-256 hash identifying an EVM code trace. Section 5.3 will describe how it is computed in our current prototype. In summary, the general form of a theorem is $\tau:=(f,\varphi(s,f.p),ph)$, which means: for every transaction invoked by calling $f$ when $\varphi$ is satisfied, if it is completed (i.e., not reverted by EVM) and its path hash equals $ph$, then all the assertions along the code path are guaranteed to hold. Obviously, this theorem is not applicable to the transaction in Attack 1 because it violates (0 $\leq$ _value < $\texttt{2}^{\texttt{255}}$) $\land$ (0 $\leq$ _fee < $\texttt{2}^{\texttt{255}}$).
The generality of the theorem can be understood in three ways. First, the guard not only constrains the referenced variables, but also implicitly states that nothing else matters, such as other parameters in the call or other storage variables. For example, addresses msg.sender, _from and _to can be arbitrarily aliased to each other, and the assertions will still hold. Second, the guard is a fairly general (i.e., weak) precondition to ensure the safety property, but it is much simpler than the weakest precondition (WP). For example, the WP will have to cover the situation $\texttt{(2}^{\texttt{255}}$ $\leq$ totalSupply < $\texttt{2}^{\texttt{256}}$), which is very complicated for integer-overflow reasoning. However, this is really unnecessary because it is hard to imagine why a cryptocurrency needs a total supply greater than $2^{255}$ (equivalent to $7*10^{66}$ per person in the world). Unlike a static verifier, TCT only requires programmers to focus on conditions within the boundary of their intentions. Third, we will explain that TCT utilizes guards to reuse theorems, which is crucial for TCT’s near-zero runtime overhead.
Theorem about the reentrant code path. We now do a similar exercise about a transaction entering via clear(...). If it does not exploit the reentrancy pattern (i.e. if transition D in Figure 3 becomes a return), then a guard like $\varphi(s,f.p)$ := (0  $\leq$  totalSupply < $\texttt{2}^{\texttt{255}}$) $\land$ (_to $\neq$ msg.sender) is sufficient to ensure the invariants. However, if the transaction exploits the reentrancy pattern as in Attack 2, the code path is different, as shown in LABEL:code:atk2. The two lines marked with “*” are repeated 9 more times. Hence the theorem for the non-reentrancy transaction is not appliable to the reentrancy one, because of the path hash mismatch. The programmer is unlikely to anticipate the reentrancy code path before the year 2016, but it is fine because security of TCT does not require the programmer to anticipate every non-intention. Since it is the attacker who discovers this unintended code path, it is the attacker’s obligation to prove the corresponding theorem. The aforementioned $\varphi(s,f.p)$ will fail the proof. The attacker can define the guard $\varphi(s,f.p)$ := (balances[msg.sender] = 0) to prove the theorem, but it makes the attack harmless as it is basically a no-op repeated 10 times.
grd: assume $\varphi(s,f.p)$;
@@
4. The TCT Protocol
@@
TCT does not make any contract-level assumption (e.g. the reference-topology, the dichotomy about “contracts of interest” vs. “environment contracts”, or claims that certain contracts have been formally verified by an auditing company). The trusted computing base (TCB) is entirely at the infrastructure level. It includes Ethereum’s existing TCB and TCT-additional TCB. The former includes the Ethereum protocol, EVM, and the verified compilation to attest a contract’s EVM code to its claimed source code (Etherscan-source-verification, 7). The latter includes the TCT protocol (in this section), the Boogie verifier, and an EVM-to-Boogie translator (in Section 5).
Ethereum $E(pre,net,s,T)$
$\triangleright$ pre-executor prepre, executor netnet, storage $s$, theorem repo $T$
type TX(f,v)(f,v) $\triangleright$ entry $f$ and values $v$ for parameters $f.p;$
$\triangleright$ Execution of a TX yields (okok, EVM-trace $ct$, path-hash $ph$).
[…]
Theorem repo $E.T$: set of Hyp
$\triangleright$ theorems of $E$, replicated everywhere
function Apply(exectrexectr: executor, $tx:$ TX, $h:$ Hyp)
  if $h.f=tx.f~\land~(tx.v,E.s)\models h.\varphi$ then
   $(ok,ct,ph)\leftarrow exectr(tx)$
   return ok&&h.ph=phok~\&\&~h.ph=ph
  end if
  return falsefalse
end function
function Execute($tx:$ TX, $h:$ Hyp)
  if $h\in E.T$ then
   $ok\leftarrow$ Apply($E.net,tx,h$)
   if okok then
     $E.net$.commitStorageUpdates()
   end if
  end if
end function
function FindTheorem($tx:$ TX)
  return ChooseOne($\{\tau\in E.T~|$~ Apply($E.pre,tx,\tau$) }\})
end function
function AddTheorem($ct:$ EVM-trace, $\varphi:$ guard)
  $ph\leftarrow E.net.computeHash(ct)$
  if $E.net.prove(\varphi\Rightarrow VC(ct)$then
   $\tau\leftarrow(ct.f,\varphi,ph)$ $\triangleright$ we have a theorem $\tau$
   $E.T\leftarrow E.T\cup\{\tau\}$
   return $\{\tau\}$
  end if
  return $\{\}$
end function
function Launch($tx:$ TX)
  $A\leftarrow$ FindTheorem(txtx) $\triangleright$ Empty if nothing is found.
  if $A\not$$\{\}$ then
   Execute(txtx, AA)
   return
  end if
  $\varphi\leftarrow$... $\triangleright$ guard comes from user
  $(ok,ct,ph)\leftarrow E.net(tx)$
  if okok then
   $B\leftarrow$ AddTheorem(ctct, $\varphi$)
   if $B\not$$\{\}$ then
     $E.net$.commitStorageUpdates()
   end if
  end if
end function
Algorithm 1 The TCT protocol.
Figure 4. (Upper) A protocol flow of adding a theorem to the repo; (Lower) A protocol flow of the repo hit scenario.
The TCT protocol piggybacks on the existing Ethereum protocol. The system is the same as the one in Figure 2, except that every node has three additional components: the EVM-to-Boogie translator, the Boogie verifier, and a theorem repository that stores all verified theorems. Note that the protocol is orthogonal to the underlying consensus mechanism, because theorem-proving is a stateless deterministic computation.
The protocol is modeled in Algorithm 1 with Figure 4 illustrating two example flows. The system with which the TX issuer interacts is $E(pre,net,s,T)$. As explained in Section 2, the consensus mechanism abstracts the whole network as a single VM, denoted as the executor netnet. The storage of netnet is $s$. A particular node in the network acts as the pre-executor prepre, which also operates on $s$. Theorem repo $T$ will be explained later. A transaction $tx$ consists of the entry function $f$ and the values $v$ for parameters $f.p$. Both netnet and prepre can execute $tx$ (denoted as $E.net(tx)$ and $E.pre(tx))$, but only netnet has the ability to commit the storage updates. An execution without a subsequent commit makes no persistent effect on $s$. Executing $tx$ results in a triple ($ok$, EVM-trace $ct$, path-hash $ph$). The Boolean $ok$ indicates whether any revert instruction is encountered. Because theorems are proven hypotheses, we refer to the type of theorem as Hyp. Theorem repository $E.T$, replicated on every node, contains all theorems that have been proven by netnet.
[…]
The TX issuer’s procedure to submit a transaction is described in Launch(tx)(tx), illustrated in Figure 4 (lower). It calls FindTheorem(tx)(tx), which consults the pre-executor prepre to find a theorem applicable to $tx$. Applicability means that (1) the theorem is about the entry function (i.e., $h.f=tx.f$), (2) the theorem’s guard is satisfied (i.e., $tx.v,E.s\models h.\varphi$), and (3) the code trace of the pre-execution matches the code trace of the theorem (i.e., $h.ph=ph$). There can be multiple applicable theorems. Function ChooseOne chooses any of them or return { } when the set is empty. Note that the pseudo code of FindTheorem shows its concise functional specification, but the actual code is very efficient: pre-execution is an existing step today, so the only additional work is to use ($f$, phph) for a dictionary lookup in $E.T$, then check $\varphi$. The first applicable theorem is returned. Once such a theorem $A$ is found, Launch(tx)(tx) calls Execute(tx,A)(tx,A) to execute the transaction by the Ethereum network netnet. We call this a repo hit. Most transactions should be repo hits, because the programmers should have tested these scenarios and pre-populated the theorems. The runtime overhead is just the applicability check, which is near zero as we will show in Section 7.4.
When there is no applicable theorem (i.e.., a repo miss), it means a real user issues a $tx$ that goes through a code path unseen in all test cases. It is likely an attack, e.g., the reentrancy attack in Section 3. However, if $tx$ is not an attack, the user needs to provide a guard $\varphi$. Launch(tx)(tx) executes $tx$ and tries to add the theorem about $\varphi$ into the repo (if the theorem is proven). The protocol flow is shown in Appendix A.2.
@@
Optimizations for low storage and communication overheads. To store theorems efficiently in the repository, all theorems that share the same entry function and guard but differ in their path hashes can be combined as a multi-path theorem, which states that, if the guard is satisfied, then any code path in this multi-path set is safe. The amortized cost to store a new code path is only one Keccak-256 value, i.e., 32 bytes. The instructions in a code path do not need to be stored, because they are only used during theorem-proving time. Once it is done, the path hash suffices. Once a theorem is stored in the repository, a Keccak-256 hash can be computed for it. We call it theorem hash. Because most transactions are repo hits, we can conclude that the network overhead of TCT is 64 bytes per transaction, due to the messages “AA” and “Execute(tx,A)Execute(tx,A)”.
5. Implementation
5.1. Overview
@@
Figure 5 shows how a theorem about a transaction is proven. We use Remix IDE (remix, 4) as the transaction issuer. The node consists of a Go-Ethereum (Geth) (geth, 20) executable and other components. From Remix, we deploy a contract and issue a transaction. Geth returns a TX-hash that identifies the transaction. Note that TX-hash is an Ethereum concept, not to be confused with our TCT concept path hash. Based on the debug_traceTransaction feature of Geth, we build a trace fetcher, which uses TX-hash to get the EVM code trace.
Figure 5. Components of the implementation.
@@
The main component we built is the concolic executor. The execution result of it is a straight-line Boogie code like the one in LABEL:code:proof. The input includes the EVM code trace, the theorem, and Solidity compilation symbol information. The properties specified in the NatSpec notation are in the symbol information. When TCT is deployed in production, the symbol information should come from a verified-compilation service, such as the Etherscan service named Verify & Publish Contract Source Code (Etherscan-source-verification, 7). The service attests that a contract’s EVM code is compiled from the claimed source code by an official Solidity compiler. Note that the NatSpec contents are contained in the symbol information.
5.2. Concolic executor
The concolic executor requires a very substantial amount of effort to implement. Due to space constraints, we can only describe it in this short subsection. It is an EVM instruction executor with only one difference from the executor in a classical EVM – the computation of each instruction is performed symbolically. The data structure that we define to hold a symbolic value is named symbolic value tree, or SVT. EVM is a stack-based machine. Unlike a classical EVM whose stack holds concrete values, our concolic executor’s stack holds SVT references. The implementation requires thorough understanding about the Solidity specification (soliditydoc, 30) and the EVM specification (wood2014ethereum, 39).